OpenStudy (anonymous):

What are the lengths of the missing sides of a 30-60-90 Degree triangle if the longer leg of the triangle is 18 centimeters. I honestly don't understand it and would like some help 

OpenStudy (wolf1728):

The sides are always in the proportion of 1 Sq Root of 3 and 2

OpenStudy (wolf1728):

I'm assuming the "longer leg" means the square root of 3 side. So the short side = 18/sq root(3)

OpenStudy (wolf1728):

And the long side is 2*(18/sq root(3))

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\left\{h=18 \sec (30 {}^{\circ})=12 \sqrt{3},a=18 \tan (30 {}^{\circ})=6 \sqrt{3}\right\} \]

OpenStudy (wolf1728):

robtobey - Those answers you gave agree with mine.

OpenStudy (anonymous):

These are the numbers I got.\[\left\{\left(12 \sqrt{3}\right)^2,\left(6 \sqrt{3}\right)^2+18^2\right\} \]\[\{432,432\} \]